# MeshSplat: Generalizable Sparse-View Surface Reconstruction via Gaussian Splatting

### v1

@@ 5 Limitations
Although our model achieves promising results in generalizable sparse-view surface reconstruction, it still has some limitations. MeshSplat sometimes may predict discontinuous depth maps in weakly textured areas, as shown in the supplimentary materials, although the rendered RGB image is reliable. This is probably due to the ambiguity of feature matching in these areas. Additionally, MeshSplat cannot reconstruct surfaces in regions that are not visible from the input views. Incorporating generative approaches might be a good choice in future work.
@@ 6 Conclusion
We present MeshSplat, a generalizable surface reconstruction framework designed for sparse-view inputs. We use 2DGS as scene representation and construct an end-to-end surface reconstruction network. To better meet the needs for more accurate prediction of 2DGS, we propose the WCD loss to regularize coarse point maps, and also use a normal prediction network to predict the orientations of 2DGS. Experimental results demonstrate that MeshSplat outperforms existing methods in generalizable sparse-view mesh extraction. Future work could explore more efficient architectures to predict the attributes of 2DGS more accurately.
@@ 7 Uncertainty-Guided Negative Log-Likelihood Loss in Normal Prediction
In this section, we explain the uncertainty-guided negative log-likelihood (UG-NLL) loss we use in the normal map supervision, which is proposed by [1].
@@ 8 Datasets
Our experiments are based on three indoor multi-view datasets, including Re10K [73], Scannet [11] and Replica [46] datasets. RE10K is licensed by Google LLC under a Creative Commons Attribution 4.0 International License. Scannet is under MIT license. Replica is under a custom license that only allows a research or educational purpose. Since Re10K does not have ground-truth meshes or point clouds, we use COLMAP [45] to reconstruct dense point clouds for 20 scenes under dense-view inputs with provided camera parameters, and take these point clouds as ground truth. Additionally, since sparse input viewpoints cannot observe the entire content of the scene, we cull ground-truth meshes or point clouds based on the input view frustums. Visualizations of some of these point clouds can be found in Figure 11.
@@ 9 Additional Implementation Details
Network Architecture. The backbone of our model follows the same structure and initialization methods as MVSplat [10]. For the cost volume construction, we sample $D=128$ depth candidates between the near and far planes. For the normal prediction network, we use a total of 10 convolutional layers to predict the 2D Gaussian orientations, and the pseudo ground-truth supervision is applied to normal maps at all scales including 1/4, 1/2 and full resolution. For the depth refinement network, they are implemented as U-Net with self-attention and cross-attention layers.
@@ 10 Additional Experimental results
MVSNeRF* [4] 0.76 0.341
SparseNeuS* [32] 7.0483 0.843
pixelSplat [3] 0.194 125.4
MVSplat [10] 0.072 12.0

### v2

@@ 5 Limitations
Although our model achieves promising results in generalizable sparse-view surface reconstruction, it still has some limitations. MeshSplat sometimes may predict discontinuous depth maps in weakly textured areas, as shown in the Appendix, although the rendered RGB image is reliable. This is probably due to the ambiguity of feature matching in these areas. Additionally, MeshSplat cannot reconstruct surfaces in regions not visible from the input views. Incorporating generative approaches might be a good choice.
@@ 6 Conclusion
We present MeshSplat, a generalizable surface reconstruction framework designed for sparse-view inputs. We use 2DGS as scene representation and construct an end-to-end surface reconstruction network. To better meet the needs for more accurate prediction of 2DGS, we propose the WCD loss to regularize coarse point maps, and also use a normal prediction network to predict the orientations of 2DGS. Experiments demonstrate that MeshSplat outperforms existing methods in generalizable sparse-view mesh extraction.
[…]
@@ 7 Uncertainty-Guided Negative Log-Likelihood Loss in Normal Prediction
In this section, we explain the uncertainty-guided negative log-likelihood (UG-NLL) loss we use in the normal map supervision, which is proposed by [bae2021estimating].
@@ 8 Datasets
Our experiments are based on three indoor multi-view datasets, including Re10K [zhou2018stereo], Scannet [dai2017scannet] and Replica [straub2019replica] datasets. RE10K is licensed by Google LLC under a Creative Commons Attribution 4.0 International License. Scannet is under MIT license. Replica is under a custom license that only allows a research or educational purpose. Since Re10K does not have ground-truth meshes or point clouds, we use COLMAP [schonberger2016colmap] to reconstruct dense point clouds for 20 scenes under dense-view inputs with provided camera parameters, and take these point clouds as ground truth. Additionally, since sparse input viewpoints cannot observe the entire content of the scene, we cull ground-truth meshes or point clouds based on the input view frustums. Visualizations of some of these point clouds can be found in Figure 11.
@@ 9 Additional Implementation Details
Network Architecture. The backbone of our model follows the same structure and initialization methods as MVSplat [chen2024mvsplat]. For the cost volume construction, we sample $D=128$ depth candidates between the near and far planes. For the normal prediction network, we use a total of 10 convolutional layers to predict the 2D Gaussian orientations, and the pseudo ground-truth supervision is applied to normal maps at all scales including 1/4, 1/2 and full resolution. For the depth refinement network, they are implemented as U-Net with self-attention and cross-attention layers.
@@ 10 Additional Experimental results
MVSNeRF* [chen2021mvsnerf] 0.76 0.341
SparseNeuS* [long2022sparseneus] 7.0483 0.843
pixelSplat [charatan2024pixelsplat] 0.194 125.4
MVSplat [chen2024mvsplat] 0.072 12.0